OpenStudy (gabbyalicorn):

A circular piece of fabric has a radius of 1.2 feet. The fabric sells for $5.40 per square foot. What is the total cost of the circular piece of fabric? A. $6.48 B. $20.35 C. $24.42 D. $40.69

OpenStudy (gabbyalicorn):

@geerky42

OpenStudy (gabbyalicorn):

@johnweldon1993 @jabez177

OpenStudy (jjuden):

its c

jabez177 (jabez177):

Direct answers Madam... She's correct though. :P

OpenStudy (jjuden):

(~3.14) * (1.2 ft)^2 = about 4.524 square feet 4.524 square feet * $5.40 = about $24.42

OpenStudy (johnweldon1993):

Correct^ The way to get square feet would be the AREA of the fabric So area of a circle = \(\large \pi r^2 = 3.14 \times (1.2)^2\) And multiply that answer by the 5.4 per square foot to get to your answer :)

OpenStudy (gabbyalicorn):

I get what you guys are saying and did the math too. :) Thx.

OpenStudy (jjuden):

you're welcome